HIGHLY ACCURATE VIBRATIONAL TRANSITION ENERGIES FOR THE HIGH ENERGY DENSITY MATERIAL TETRAHEDRAL N_{4}

\underline{TIMOTHY J. LEE}, MS T27B-1, NASA Ames Research Center, Moffett Field, CA 94035-1000; JAN M.L. MARTIN, Department of Organic Chemistry, Weizmann Institute of Science, 76100 Rehovot, Israel.

More than a decade ago, tetrahedral (T_{d}) N_{4} was suggested as a potential high energy density matrial, which could be used as a novel rocket fuel, among other possible uses. However, the only report of its possible synthesis was inconclusive, partly because the isotopic shifts found in the matrix isolation experiment did not agree well with theoretical values obtained from harmonic frequencies. In this study we have re-examined the theoretical spectroscopy of T_{d} N_{4} using very high levels of theory and by computing a full quartic force field. Vibrational frequencies obtained from this new force field, together with isotopic shifts, will be reported. As a by product of this study, we have also computed a new more accurate heat of formation for T_{d} N_{4} which confirms its potential as a high energy density material.